OpenStudy (azionne):

5(x - 4) = 12 will medal

OpenStudy (calculusxy):

\[\large 5(x - 4) = 12 \rightarrow 5(x) + 5(-4) = 12 \]

OpenStudy (pawanyadav):

5x-20=12 5x=20+12=32 So, divided by 5 both side x=32/5

OpenStudy (calculusxy):

I think that you should be able to solve from there. After all, @pooja195 has been helping you quite a lot with these questions as far as I've seen :)

OpenStudy (azionne):

im still rusty

OpenStudy (azionne):

@pooja195

rvc (rvc):

multiply 5(x-4)=12

OpenStudy (azionne):

5x-20=12

OpenStudy (azionne):

5x=32

rvc (rvc):

now u need to isolate x so add 20 on both sides

rvc (rvc):

correct u got it :)

OpenStudy (azionne):

x=0.15625

OpenStudy (azionne):

but what is this answer in fraction form

OpenStudy (azionne):

the problem is asking for a fraction

rvc (rvc):

this is in decimal form the answer in fraction form would be : x=32/5

OpenStudy (azionne):

its telling me to answer in mixed numbers

rvc (rvc):

so convert it into mixed fraction

pooja195 (pooja195):

Not mixed just a fraction x=32/5

OpenStudy (azionne):

its asking for mixed numbers